I am being asked to write an equation of a line perpendicular to the given line. the equation is x -4y =20. I am not sure how to do this.
Answer by nerdybill(2448)

I am being asked to write an equation of a line perpendicular to the given line. the equation is x -4y =20
.
First, determine the slope of:
x -4y =20
Do this by manipulating it so it is in the slope-intercept form:
y = mx + b
where
m is the slope
.
x -4y =20
-4y = -x + 20
Dividing both sides by -4:
y = (1/4)x - 5
We see that it has a slope of 1/4
.
If a line is perpendicular, the slopes are negative reciprocal:
If our new slope is m:
m(1/4) = -1
m = -4 (our new slope)
.
So, all we need to do is change the slope of:
y = (1/4)x - 5
to -4
y = -4x - 5 (this is what they are looking for)